the sum of two numbers is twenty-five. the larger number is seven less than four times the smaller number. find the numbers.
Answer by mananth(16949)

let one number be x
the other number will be 25-x
..
25-x =4x-7
4x+x=25+7
5x=32
x= 6.4
The other number will be 18.6